Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Message ID: 243     Entry time: Thu Mar 28 10:34:41 2019
Author: Drive Team- Ino&Thierry&Armand + Juli (IFAE) 
Type: Telescope Operations 
Category:  
Subject: High speed drive tests 
Status:  

new session of high speed drive tests after issues with Power Supply.

  • Tests of emergency stop (with a speed at 25% and 43 % GRB speed) with a modification of the jerk value from 0.08 to 0.2 ==> OK
  • Tests of azimuth movements : (-90 to 270 degres)
    • at 60% of GRB speed. ==> Ok for 6 movements
    • at 80% of GRB speed
      • Ok for 4 movements
      • Failures due to resistance  (sur-tension). Breaking modules in failure then reset
      • Ok for 2 next movements
    • at 100% of GRB speed :
      • 3 movements from -90 to 90 ==> OK
      • Temperature of breaking resistance are getting very high (at least the top one) 
ELOG V3.1.5-fc6679b